# Balancing Confidentiality and Transparency for Blockchain-based Process-Aware Information Systems

### v1

@@ 4.2 Information viewpoint
The inspect functionality pertains to the access to message data. The public data can be freely accessed, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above). To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester. This is the case, e.g., for the accepted field of the “Check availability” task. Reading confidential data requires more passages, illustrated in Fig. 6, and requires a preliminary step. All users should request at least once (and at every attribute update) a decryption key (Fig. 6(a)), to be used for every message they are going to (Fig. 6). For the first one, namely the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a). The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger. Then, (4.3a) the Confidentiality Manager generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via Attribute-Based Encryption (ABE), and (4.4a) sends it back. The decryption of confidential data is depicted in Fig. 6(b). Once the process participant has received the a-b key, they can use it to decrypt the confidential data. To do so, (4.0b) the user makes a request to the Confidentiality Manager. Then, (4.1b) the Confidentiality Manager retrieves the encrypted data from the Distributed FS and (4.2b) forwards it to the user. At this point, the latter (4.3b) uses the a-b key to decrypt the confidential data via ABE. Only if the user’s attributes satisfy the policy previously used to encrypt the data, can the user’s a-b key decrypt the confidential data (and the user inspect it in-clear).
@@
The a-b key of user 0xc5…Fd43 (PID476948 and PATIENT) can decrypt the report instead, but not the other two documents.
@@ 5 Evaluation
5.1 CONFETTY Implementation and Evaluation
Figure 7 illustrates the technologies utilized in the implementation of CONFETTY. Our prototype is deployed on three main tiers. An application server hosts the Confidentiality Manager and the Process Manager, developed in Python 3.6.9 and Java SDK-13, respectively. Users communicate with it via TLS communication channels [37]. We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache. Accessed: December 7, 2024. as the Distributed Tamper-proof File Storage. We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js.1 The source code of CONFETTY, alongside the experimental data we use, is publicly available at github.com/apwbs/CONFETTY.
@@ 5.2 Threat Model
The CONFETTY boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditory). In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented. As for tampering, trust is based on the fairness of the user who sends the data. It cannot be the case for a user to repudiate an action because all of them are recorded on the blockchain with a transaction signed with the user’s private key. An attacker could sniff the data a user sends to CONFETTY. However, for those communications we resort to TLS channels, specifically designed to prevent information leakage [37]. DoS attacks are in principle feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform. Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1.0 in Sect. 4.2) which malicious nodes could exploit with multiple calls and bulky payloads. However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss. Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE. Elevating a user’s privileges is not possible since attributes and permissions are linked to a blockchain account.
The Blockchain Platform is the trusted component in the Blockchain Boundary (BB). In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the real one. This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their own private keys. Information disclosure can be achieved only on public data or resource locators of sensitive data. However, the latter is stored upon ABE-encryption, making its content unreadable by malicious users. Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols.
As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack, and elevation of privileges, are counteracted by design, too. Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain. Since we resort to ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.

### v2

@@ 4.2 Information viewpoint
The inspect functionality pertains to the access to message data. The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above). To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester. This is the case, e.g., for the accepted field of the “Check availability” task. Reading confidential data requires more passages, illustrated in Fig. 6, and requires a preliminary step. All users should request at least once (and at every attribute update) a decryption key (Fig. 6(a)), to be used for every message they are going to read (Fig. 6(b)). For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a). The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger. Then, (4.3a) the Confidentiality Manager generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via Attribute-Based Encryption (ABE), and (4.4a) sends it back. The decryption of confidential data is depicted in Fig. 6(b). Once the process participant has received the a-b key, they can use it to decrypt the confidential data. To do so, (4.0b) the user makes a request to the Confidentiality Manager. Then, (4.1b) the Confidentiality Manager retrieves the encrypted data from the Distributed FS and (4.2b) forwards it to the user. At this point, the latter (4.3b) uses the a-b key to decrypt the confidential data via ABE. Only if the user’s attributes satisfy the policy previously used to encrypt the data, can the user’s a-b key decrypt the confidential data (and the user inspect it in-clear).
@@
The a-b key of user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
@@ 5 Evaluation
5.1 Implementation
Figure 7 illustrates the technologies utilized in the implementation of CONFETTY. Our prototype is deployed on three main tiers. An application server hosts the Confidentiality Manager and the Process Manager, developed in Python 3.6.9 and Java SDK-13, respectively. Users communicate with it via TLS communication channels [37]. We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache. Accessed: December 13, 2024. as the Distributed Tamper-proof File Storage. We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js.1 The source code of CONFETTY, alongside the experimental data we use, is publicly available at github.com/apwbs/CONFETTY.
@@ 5.2 Threat Model
The CONFETTY Boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditor). In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented. As for tampering, trust is based on the fairness of the user who sends the data. It cannot be the case for a user to repudiate an action because all of them are recorded on the blockchain with a transaction signed with the user’s private key. An attacker could sniff the data a user sends to CONFETTY. However, for those communications we resort to TLS channels, specifically designed to prevent information leakage [37]. DoS attacks are in principle feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform. Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1.0 in Sect. 4.2) which malicious nodes could exploit with multiple calls and bulky payloads. However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss. Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE. Elevating a user’s privileges is not possible since attributes and permissions are linked to a blockchain account.
The Blockchain Platform is the trusted component in the Blockchain Boundary (BB). In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the real one. This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their own private keys. Information disclosure can be achieved only on public data or resource locators of sensitive data. However, the latter is stored upon ABE-encryption, making its content unreadable by malicious users. Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols. As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack, and elevation of privileges, are counteracted by design, too. Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain. Since we resort to ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.